A survey questioned 1000 people regarding raising the legal drinking age from 18 to 21.
Of the 560 who favored raising the age, 390 were female. Of the 440 opposition
responses, 160 were female. If a person selected at random from this group is a man,
what is the probability that the person favors raising the drinking age?
1
Expert's answer
2021-09-14T06:06:03-0400

SOLUTION


The probability that the person favors raising the drinking age is given as:

"Pr\left(Favors\:raising\:the\:drinking\:age/\:Male\:\right)=\frac{Total\:number\:of\:males\:who\:favors\:drinking\:age}{Total\:Number\:of\:Males}"


"Pr\left(Favors\:raising\:the\:drinking\:age/\:Male\:\right)=\frac{\left(560-390\right)}{\left(560-390\right)+\left(440-160\right)}"


"Pr\left(Favors\:raising\:the\:drinking\:age/\:Male\:\right)=\frac{170}{170+280}=\frac{170}{450}=\frac{17}{45}"


Hence, the probability that the person favors raising the drinking age is "=\frac{17}{45}"